If c is the harmonic mean between a and b, then (c ⁄ a) + (c ⁄ b) is equal to

Grade:12

1 Answers

bharat bajaj IIT Delhi
9 years ago
c = \frac{2ab}{a+b}
\frac{c}{a}= \frac{2b}{a+b}
\frac{c}{b} = \frac{2a}{a+b}
\frac{c}{a} + \frac{c}{b} = 2
Thanks&Regards
Bharat Bajaj
askiitian faculty
IIT Delhi